The jet stream is a strong wind current that flows across the United States. Flying with the jet stream, a plane flew 1872 miles in 3 hours. Against the same wind, the trip took 4 hours. Find the speed of the plane in still air and the speed of the wind current.
Answer by lwsshak3(11628)

The jet stream is a strong wind current that flows across the United States. Flying with the jet stream, a plane flew 1872 miles in 3 hours. Against the same wind, the trip took 4 hours. Find the speed of the plane in still air and the speed of the wind current
**
let x=speed of plane in still air
let c=speed of wind current
Travel time=distance/speed
Speed of plane with the wind=x+c
Speed of plane against the wind=x-c
..
1872/(x+c)=3
1872/(x-c)=4
..
1872=3x+3c
1872=4x-4c
..
7488=12x+12c (multiply by 4)
5616=12x-12c (multiply by 3)
add
13104=24x
x=546 mph
..
subtract
1872=24c
c=78 mph
ans:
speed of plane in still air=546 mph
speed of wind current=78 mph